OpenStudy (anonymous):

Two points sit on a line with the following coordinates: P (3, 3) and Q (4, 8). Which is the distance from P to Q? A. sqrt 10 B. sqrt 15 C. sqrt 23 D. sqrt 26

ganeshie8 (ganeshie8):

use the distance formula

ganeshie8 (ganeshie8):

distance = \(\large \sqrt{(x2-x1)^2 + (y2-y1)^2}\)

OpenStudy (anonymous):

Say that x1=3,y1=1,x2=4,y2=4 (the coordinates of the points). Then the distance will be: \[d =\sqrt{(y2-y1)^2+(x2-x1)^2}\]

OpenStudy (johnweldon1993):

\[\sqrt{(y_2 - y_1)^2 + (x_2 - x_1)^2 }\] Your 2 points are P (3, 3) and Q (4, 8). (x1,y1) (x2,y2)

OpenStudy (nirmalnema):

root over (4-3)^2+(8-3)^2 root over 1^2+5^2 =sq rt of26